Two resistors, 36.0 n and 78.0 n, are connected in parallel. The current through the second resistor is 2.70 A. (a) Determine the current in the first resistor? (b) What is the total power supplied to the two resistors?
Answer #1

current through second resistor = I*R1 / (R1 + R2)

2.7 = I*36 / (36 + 78)

I = 8.55 A

(a) current in the first resistor = I*R2 / (R1 + R2) = 8.55 *78 / (36 + 78) = 5.85 A

(b) total current = 5.85 + 8.55 = 14.4 A

total resistance = R1R2/(R1 +R2) = 36*78/ (36 + 78) = 24.63 ohm

TOTAL POWER = I^2*R = 14.4^2*24.63 = 5107.6 W